# Differentially-private frugal estimation of quantiles

### v1

@@ IX Experimental Results
In this section we present and discuss the results of the experiments carried out for Frugal-1U-L, Frugal-1U-G, Frugal-1U-$\rho$, Frugal-2U-SA and LDPQ. The source code has been compiled using the Apple clang compiler v15.0 with the following flags: -Os -std=c++14 (it is worth recalling that on macOS the flag -Os optimizes for size and usually on this platform the resulting executable is faster than the executable obtained by compiling using the -O3 flag). The tests have been carried out on an Apple MacBook Pro laptop equipped with 64 GB of RAM and a 2.3 GHz 8-core Intel Core i9. The experiments have been repeated ten times for each specific category of test and the results have been averaged; the seeds used to initialize the pseudo-random generators are the same for each experiment and algorithm being tested.
The tests have been performed on eight synthetic datasets, whose properties are summarized in Table I. The experiments have been executed varying the stream length, the quantile, the privacy budget, either $\epsilon$ or $r$ (which represents the LDPQ privacy budget), $\delta$ and $\rho$. Table II reports the default settings for the parameters.
We compare Frugal-1U-L with LDPQ. In particular, we remark here that this is possible owing to the fact that the values selected for $r$ correspond exactly to the values selected for $\epsilon$, since in [51] the authors report that their algorithm satisfies $\epsilon$-LDP with $\epsilon=\ln((1+r)/(1-r))$ (alternatively, $r=(e^{\epsilon}-1)/(e^{\epsilon}+1)=\tanh(\epsilon/2))$). Therefore, the provided comparison between Frugal-1U-L and LDPQ is fair.
Additionally, we report here the results obtained for Frugal-1U-G and Frugal-1U-$\rho$. Moreover, we also report the result for Frugal-2U-SA.
TABLE II: Default settings of the parameters.
@@
chunks $\{2,4,8,16\}$ 4
(a) privacy budget: $\epsilon$ and $r$
@@
Figure 6: Frugal-1U-L versus LDPQ. Relative error varying the privacy budget $\epsilon$ and $r$, the quantile $q$ and the stream size $n$. Throughput measured in updates/s.
@@
Figure 9: Frugal-1U-L versus LDPQ: relative error varying the distributions.
(a) $\rho$
(b) Quantile
(c) $n$
(d) Chunks
Figure 10: Frugal-2U-SA. Relative error varying the privacy budget $\epsilon$, the quantile $q$, the stream size $n$ and the number of chunks.
Figure 11: Frugal-2U-SA. Relative error varying the distributions.
@@
The experimental results for Frugal-1U-L (using the Laplace mechanism) versus LDPQ are shown in Figure 6. Regarding Frugal-1U-L, as depicted in Figure 6(a), the relative error decreases as expected when the privacy budget $\epsilon$ increases, meaning that the utility (see Section II) of the released value increases when $\epsilon$ increases. Therefore, a good tradeoff between privacy and utility is reached for $0.5\leq\epsilon\leq 1$. On the contrary, LDPQ presents a very high relative error between 10 and 100, making the output useless from a practical perspective.
Figure 6(b) depicts the relative error varying the computed quantile. The observed trend is the same, with Frugal-1U-L outperforming LDPQ. As shown in Figure 6(c), the stream size does not affect the security of the released quantile in the case of Frugal-1U-L whilst LDPQ exhibits a fluctuating behaviour with a relative error between 10 and 100. Finally, Figure 6(d) depicts the throughput measured in updates/s. As shown, Frugal-1U-L is up to 7 times faster than LDPQ, making it suitable for processing high speed streams.
@@
Regarding Frugal-1U-$\rho$, Figure 8(a) shows that, as expected, increasing the privacy budget $\rho$ the relative error decreases and correspondingly the utility increases. A good tradeoff between privacy and utility is reached for $0.5\leq\rho\leq 1$. Figure 8(b) and 8(c), related respectively to the relative error varying the computed quantile and the stream size present the same behaviour illustrated for the Gaussian mechanism. This is not surprising, since this mechanism adds Gaussian noise (even though the way noise is derived is of course different).
We now turn our attention to what happens when we vary the distribution. Figure 9 provides the results for Frugal-1U-L, Frugal-1U-G, Frugal-1U-$\rho$ and LDPQ. As shown, the relative error between the true quantile and the DP quantile estimate released by one of the algorithms varies with the distribution. However, for our proposed algorithms, as expected (since the global sensitivity is just 2) the algorithms can be used independently of the actual distribution, with the notable exception related to the Cauchy distribution (which can be considered adversarial for our algorithms based on Frugal-1U as discussed in [32]). LDPQ, as shown, exhibits a relative error varying between 1 and more than 100, making it useless for all of the distributions under test.
Our results show that, having fixed a distribution, the behaviour of our algorithms based on Frugal-1U does not depend on the seed used to initialize the pseudo-random number generator used to draw samples from the distribution. In this sense, our algorithms are robust. On the contrary, LDPQ heavily depends on the seeds used, exhibiting a rather large variance in the obtained results.
@@
Frugal-2U-SA results are shown in Figure 10, with regard to varying the privacy budget $\epsilon$ (Figure 10(a)), the quantile $q$ (Figure 10(b)), the stream size $n$ (Figure 10(c)) and the number of chunks used in the Sample and Aggregate framework (Figure 10(d)). The results obtained are, apparently, similar to those related to Frugal-1U-L, Frugal-1U-G and Frugal-1U-$\rho$. Increasing the chunks we observe less relative error, as expected (but we need to recall here that, in turn, this also requires a suitable chunk size in order for the algorithm to converge to a correct estimate in each chunk). Even though the obtained results are good enough, we need to take into account that these results are related to the normal distribution, and we also need to take into account the distribution parameters, which determine the minimum and maximum value used in the Sample and Aggregate framework.
Indeed, since in this case the global sensitivity is unbounded, we cannot hope to get consistently good results independently of the distribution. We confirm this in Figure 11, in which we show that the relative error associated to the release of the estimate for the 0.99 quantile using a privacy budget $\epsilon=1$ and 4 chunks may be, depending on the selected distribution, orders of magnitude higher than acceptable values that lead to a good tradeoff between privacy and utility.
As a consequence, despite the faster convergence speed to the quantile estimate, Frugal-2U-SA cannot be used to ensure privacy of the released quantile without being fully aware of the implications related to the distribution, and in particular the range (i.e., the difference between the maximum and the minimum value). Therefore, we recommend using Frugal-1U-L, Frugal-1U-G and Frugal-1U-$\rho$ instead, since these does not require distributional assumptions, are fast and require just one unit of memory.
Regarding the accuracy of Frugal-2U-SA, it can be easily derived taking into account that the noise added to the answer is $\operatorname{Lap}(\frac{u-l}{k\epsilon})$; thus, Frugal-2U-SA is $(\alpha,\beta)$-accurate with
$\alpha=\frac{\ln\left(\frac{1}{\beta}\right)(u-l)}{k\epsilon}.$ (11)
Therefore, fixing the values of $\beta$ and $\epsilon$, $\alpha$ depends both on the range $u-l$ (i.e., it depends on the actual distribution from which the stream items are drawn) and on $k$ (i.e., the number of chunks in which the stream is partitioned). In order to improve the accuracy, it is consequently desirable to use a number of chunks as large as possible (ideally, $k=u-l$). As we have already noted, in order to increase the value of $k$ we need, however, a sufficiently large stream size.
X Conclusions
In this paper, we proposed DP algorithms for tracking quantiles in a streaming setting. Our algorithms are DP variants of the well-known Frugal-1U and Frugal-2U algorithms, characterized by the property of requiring just a tiny amount of memory to process a stream while guaranteeing surprising accuracy for the estimates of a quantile. In particular, for Frugal-1U we gave corresponding $\epsilon$-DP, $(\epsilon,\delta)$-DP, and $\rho$-zCDF algorithms after proving that the global sensitivity of Frugal-1U is equal to 2. Moreover, we compared our Frugal-1U-L algorithm with LDPQ, which is a recent state of the art algorithm, and showed that our algorithm outperforms LDPQ with regard to both accuracy and speed.
For Frugal-2U, we showed that the global sensitivity is unbounded, by providing a carefully crafted adversarial stream, which was also analyzed from a probabilistic perspective. Then, we provided an $\epsilon$-DP algorithm based on the Sample and Aggregate framework. The algorithms are simple to implement; however, due to the potential issues related to Frugal-2U (unbounded global sensitivity and dependency on the distribution range) we recommend the use of Frugal-1U-L, Frugal-1U-G, Frugal-1U-$\rho$, which exhibit good accuracy and security, as shown in the extensive experimental results provided.

### v2

@@ 5 Experimental Results
In this section we present and discuss the results of the experiments carried out for Frugal-1U-L, Frugal-1U-G, Frugal-1U-$\rho$. The experiments are limited to our algorithms owing to the fact that, to the best of our knowledge, there are no other frugal algorithms for the estimation of quantiles designed for the central model of differential privacy. The source code has been compiled using the Apple clang compiler v15.0 with the following flags: -Os -std=c++14 (it is worth recalling that on macOS the flag -Os optimizes for size and usually on this platform the resulting executable is faster than the executable obtained by compiling using the -O3 flag). The tests have been carried out on an Apple MacBook Pro laptop equipped with 64 GB of RAM and a 2.3 GHz 8-core Intel Core i9. The experiments have been repeated ten times for each specific category of test and the results have been averaged; the seeds used to initialize the pseudo-random generators are the same for each experiment and algorithm being tested.
The tests have been performed on eight synthetic datasets, whose properties are summarized in Table 1. The experiments have been executed varying the stream length, the quantile, the privacy budget, $\epsilon$, $\delta$ and $\rho$. Table 2 reports the default settings for the parameters.
Table 2: Default settings of the parameters.
@@
(a) privacy budget
@@
Figure 1: Frugal-1U-L. Relative error varying the privacy budget $\epsilon$, the quantile $q$ and the stream size $n$. Throughput measured in updates/s.
@@
Figure 4: Frugal-1U-L: relative error varying the distributions.
@@
The experimental results for Frugal-1U-L (using the Laplace mechanism) are shown in Figure 1. As depicted in Figure 1(a), the relative error decreases as expected when the privacy budget $\epsilon$ increases, meaning that the utility (see Section 2) of the released value increases when $\epsilon$ increases. Therefore, a good tradeoff between privacy and utility is reached for $0.5\leq\epsilon\leq 1$.
Figure 1(b) and Figure 1(c) depict the relative error varying respectively the computed quantile and the stream size, showing that the two quantities do not affect the security of the released quantile. Finally, Figure 1(d) depicts the throughput measured in updates/s.
@@
Regarding Frugal-1U-$\rho$, Figure 3(a) shows that, as expected, increasing the privacy budget $\rho$ the relative error decreases and correspondingly the utility increases. A good tradeoff between privacy and utility is reached for $0.5\leq\rho\leq 1$. Figure 3(b) and 3(c), related respectively to the relative error varying the computed quantile and the stream size present the same behaviour illustrated for the Gaussian mechanism. This is not surprising, since this mechanism adds Gaussian noise (though the way noise is derived is of course different).
We now turn our attention to what happens when we vary the distribution. Figure 4 provides the results for Frugal-1U-L, Frugal-1U-G and Frugal-1U-$\rho$. As shown, the relative error between the true quantile and the DP quantile estimate released by one of the algorithms varies with the distribution. However, for our proposed algorithms, as expected (since the global sensitivity is just 2) the algorithms can be used independently of the actual distribution, with the notable exception related to the Cauchy distribution (which can be considered adversarial for our algorithms based on Frugal-1U as discussed in [32]).
Our results show that, having fixed a distribution, the behaviour of our algorithms based on Frugal-1U does not depend on the seed used to initialize the pseudo-random number generator used to draw samples from the distribution. In this sense, our algorithms are robust.
@@
6 Conclusions
In this paper, we proposed DP algorithms for tracking quantiles in a streaming setting. Our algorithms are DP variants of the well-known Frugal-1Ualgorithm, characterized by the property of requiring just a tiny amount of memory to process a stream while guaranteeing surprising accuracy for the estimates of a quantile. In particular, for Frugal-1U we gave corresponding $\epsilon$-DP, $(\epsilon,\delta)$-DP, and $\rho$-zCDF algorithms after proving that the global sensitivity of Frugal-1U is equal to 2. Finally, we also showed that the proposed algorithms achieve good accuracy in the experimental results.
Declaration on Generative AI
The author(s) have not employed any Generative AI tools.